# DiffSSC: Semantic LiDAR Scan Completion using Denoising Diffusion Probabilistic Models

### v1

@@ III Methodology
Given a raw LiDAR point cloud, our objective is to estimate a more complete semantic point cloud, including unobserved points with associated semantic labels within gaps and occluded regions. As illustrated in the Fig. 2, we achieve this through a diffusion model as the core component, supported by a pre-processing semantic segmentation module and a post-processing refinement module. First, the raw LiDAR point cloud is semantically segmented using a Cylinder3D [7] to generate initial semantic logits. Next, we upsample the semantic point cloud to increase point density for the diffusion process. The duplicated semantic points undergo a forward diffusion and a reverse denoising process to adjust their positions and semantics. Notably, the semantic point cloud also serves as a conditional input for the diffusion model, guiding the generation process. The generated scene includes semantic points located in gaps and occluded areas. To further enhance the quality of the generated scene, we designed a refinement model based on MinkUnet [24, 25, 26, 27] to densify the point cloud.
@@ III-B Diffusion Semantic Scene Completion
As mentioned in Sec. I, by learning scene priors, the model gains the ability to estimate a complete scene (posterior) from partial observations (likelihood). The diffusion model efficiently learns the distribution of the ground truth data, acquiring knowledge of the scene prior. To achieve this, we gradually add noise to the ground truth $\mathcal{Y}$, resulting in $\mathcal{Y}_{1},\dots,\mathcal{Y}_{T}$, until $\mathcal{Y}_{T}$ approximates a Gaussian distribution. This form of data augmentation can be simplified using Eq. 1. However, this approach does not directly apply to semantic scene completion. In Eq. 1, the diffusion process is defined as a combined distribution of the sample $\bm{x}_{0}$ and global noise $\bm{\epsilon}\sim\mathcal{N}(0,I)$, with coefficients $\sqrt{\bar{\alpha}_{t}}$ and $\sqrt{1-\bar{\alpha}_{t}}$ controlling the ratio of noise and sample at different time steps. This mechanism was originally designed for images, which are flattened and normalized before global noise is added. Similarly, this approach has also shown good results in shape completion, as the shapes are generally close to a 3D Gaussian distribution. However, in large-scale scenes, the data distribution deviates significantly from a Gaussian, particularly due to varying data ranges across different point cloud axes. Applying global noise to the entire scene point cloud as a single entity can obscure important details, failing to account for the unique characteristics and distribution of each point. Therefore, we reformulate the diffusion process as a noise offset $\bm{\epsilon}\in\mathbb{R}^{3+C}$ added locally to each point $\bm{y}^{m}\in\mathcal{Y}$, as shown in the following equation:
@@ III-C Denoiser Design and Regularization
Since noise is added to local points during the diffusion process, we employ the commonly used $L_{2}$ loss between the added noise and the model’s predictions. Additionally, we design a regularization term for the global characteristics of the noise, by specifically focusing on the mean $\mu_{\epsilon}$ and variance $\sigma_{\epsilon}$. Thus, the loss can be formulated as follows:

### v2

@@ III Methodology
Given a raw LiDAR point cloud, our objective is to estimate a more complete semantic point cloud, including unobserved points with associated semantic labels within gaps and occluded regions. As illustrated in the Fig. 2, we build a diffusion model based on LiDiff [23], supported by a semantic segmentation module and a refinement module. First, the raw LiDAR point cloud is semantically segmented using a Cylinder3D [7] to generate initial semantic logits. Next, we upsample the semantic point cloud to increase point density for the diffusion process. The duplicated semantic points undergo a forward diffusion and a reverse denoising process to adjust their positions and semantics. Notably, the semantic point cloud also serves as a conditional input for the diffusion model, guiding the generation process. The generated scene includes semantic points located in gaps and occluded areas. To further enhance the quality of the generated scene, we designed a refinement model based on MinkUnet [24, 25, 26, 27] to densify the point cloud.
@@ III-B Diffusion Semantic Scene Completion
As mentioned in Sec. I, by learning scene priors, the model gains the ability to estimate a complete scene (posterior) from partial observations (likelihood). The diffusion model efficiently learns the distribution of the ground truth data, acquiring knowledge of the scene prior. To achieve this, we gradually add noise to the ground truth $\mathcal{Y}$, resulting in $\mathcal{Y}_{1},\dots,\mathcal{Y}_{T}$, until $\mathcal{Y}_{T}$ approximates a Gaussian distribution. This form of data augmentation can be simplified using Eq. 1. However, this approach does not directly apply to semantic scene completion. In Eq. 1, the diffusion process is defined as a combined distribution of the sample $\bm{x}_{0}$ and global noise $\bm{\epsilon}\sim\mathcal{N}(0,I)$, with coefficients $\sqrt{\bar{\alpha}_{t}}$ and $\sqrt{1-\bar{\alpha}_{t}}$ controlling the ratio of noise and sample at different time steps. This mechanism works well in shape completion, as the shapes generally approximate a 3D Gaussian distribution. However, the data distribution in large-scale scenes deviates significantly from Gaussianity, particularly due to varying data ranges across different axes. Applying global noise to the entire scene point cloud as a single entity can obscure important details. Therefore, we use the local point diffusion proposed by Nunes et al. [23], which reformulates the diffusion process as a noise offset $\bm{\epsilon}\in\mathbb{R}^{3+C}$ added locally to each point $\bm{y}^{m}\in\mathcal{Y}$, as shown in the following equation:
@@ III-C Denoiser Design and Regularization
As proposed by Nunes et al. [23], the regularization loss consists of a local term focusing on each point and a global term addressing the statistical characteristics. We extend this design to the semantic domain by employing an $L_{2}$ loss between the added noise and the model’s predictions, along with losses for the mean $\mu_{\epsilon}$ and variance $\sigma_{\epsilon}$ to regulate the global statistical properties. Thus, the overall loss is formulated as follows: